OpenStudy (anonymous):

expand the logarithmic expression log(x-5)+ log(x+2)

OpenStudy (anonymous):

using one of logarithm's laws: log(x-5)+log(x+2) = log(x-5)(x+2) = log(x^2-3x-10)

OpenStudy (anonymous):

is that it?

OpenStudy (anonymous):

yes